OpenStudy (anonymous):

how do you solve this equation: 3y+2x+12=0 (slope intercept form)

OpenStudy (anonymous):

Slope intercept form means an equation in the form y=mx + b. so, you need to rearrange the equation to be in this form. Do you know how to start that?

OpenStudy (anonymous):

i understand if it is like: 2x-y=-1 but not when there is 2 addition symbols.

OpenStudy (anonymous):

It is the same principle. You can still subtract the same number from both sides. So, you can subtract 12 from both sides -> 3y+2x=-12

OpenStudy (anonymous):

is the final answer y=2/3x+4?

OpenStudy (anonymous):

that is close, but there are some (-) signs missing.

OpenStudy (anonymous):

so y=-2/3x-4?

OpenStudy (anonymous):

Yep - that looks good! :)

OpenStudy (anonymous):

thank you